# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ III.7 Invertibility of the quantum Fisher information matrix
Beyond the issues discussed above, the invertibility of the Fisher information matrices plays a crucial role in multiparameter estimation. In the case of an invertible QFIM but singular (non-invertible) CFIM, the issue is associated to the incompleteness of the chosen measurement setup. Therefore, the solution is to enlarge the measurement space—for instance, by replacing projective measurements with generalized POVMs, or by employing sequential projective measurements [116, 117, 118]. On the other hand, a singular QFIM is more problematic as it directly implies that the CFIM is necessarily singular for all possible choices of the measurement setup. Since the QCRB is only well defined for an invertible QFIM, we strictly require
@@
When $\text{det}[\hat{\mathcal{I}}]=0$, the QCRB becomes undefined and nothing meaningful can be said about the achievable precision for the parameter set $\vec{\omega}$ [119, 120, 121, 122, 116, 66, 123]. A singular QFIM implies diverging variances and thus a vanishing SNR for parameter estimation. Such singularities typically arise from underlying dependencies between the original parameter set, and are such that signals corresponding to different parameters are indistinguishable in the measurement data.
@@ III.8 From abstract to applied
To provide a practical intuition, in Section IV we revisit some of the core concepts of parameter estimation within the context of a concrete physical example: Ramsey interferometry. This paradigmatic quantum sensing protocol provides a natural setting to illustrate how quantities like the SNR, QFI and the CRB arise in practice. We will further interrogate how the choice of measurement and different quantum resources—such as entanglement and squeezing—can be utilized to profoundly impact the ensuing sensitivity to designated parameters to be inferred $\omega$. This allows us to introduce the fundamental precision limits of classical and quantum sensing: the standard quantum limit, and the Heisenberg limit. By working through this example step by step, we reinforce the abstract ideas introduced here and lay the groundwork for more advanced schemes discussed later in this tutorial.
Building on the paradigmatic sensing strategy introduced in Sec IV, we turn to the harmonic oscillator and Landau–Zener models as minimal settings to explore quantum criticality in metrology in Sec V and VI respectively. These models allow us to highlight the defining features that make critical systems powerful platforms for sensing, such as their geometric structure and diverging susceptibilities. We present a detailed analysis of both single- and multiparameter estimation, focusing on the fundamental precision trade-offs, measurement incompatibility, and the emergence of singularities in the QFIM and how to deal with them. Particular focus is devoted to the role of resources, such as interrogation time, and to the impact of thermal fluctuations on critical quantum sensing.
@@ IV Parameter estimation in quantum metrology on the example of Ramsey interferometry
Importantly, QFI can be thought of as the SNR optimized over all possible choices of measurement. However, this optimal choice may not be directly implementable. Nonetheless, it provides a crucial benchmark: by establishing the ultimate precision limit, it allows us to assess how well feasible measurement strategies perform in comparison, and to understand how close we are to the fundamental quantum bound. Consequently, a central concept in metrology and parameter estimation is the SNR—quantifying the quality of a specific measurement by comparing the strength of the signal to the noise that obscures it. With the QFI playing the role of an optimal SNR, this establishes the ultimate benchmark for precision and the link between abstract theoretical limits, to real-world measurement challenges.
@@ IV.1 Ramsey interferometry
To illustrate the principles of quantum metrology, we consider the Ramsey interferometer—a cornerstone of precision measurement and the operating principle behind atomic clocks, which rank among the most accurate instruments ever built [124]. Their extraordinary precision stems from the inherent stability of atomic transition frequencies: each atom functions as an almost perfect clock, its ticks dictated by fundamental physical laws, much like those of an ideal pendulum. By counting the oscillations of an atom’s quantum state, time can be tracked with remarkable accuracy. In practice, however, imperfections such as fluctuations in the probing laser’s intensity and frequency introduce unavoidable sources of imprecision.
@@
In a Ramsey interferometer (see Fig. 2 for a schematic Bloch sphere representation), parameter estimation begins by preparing a quantum system in a well-defined initial state, typically the collective ground state. The first step is to apply a pulse that creates a coherent superposition of the two energy levels, represented as a quantum state vector on the Bloch sphere, halfway between the two energy eigenstates.
Next, the system evolves freely under the influence of the unknown parameter, such as a frequency, magnetic or gravitational field. During this evolution, a relative phase accumulates between the components of the superposition state, encoding information about the parameter. The evolution time is carefully chosen to maximize sensitivity to small variations in the parameter. Because directly measuring the phase is challenging, a second pulse is applied after the evolution period. This pulse converts the accumulated phase information into a measurable quantity, typically the population difference between the two energy levels. The population difference depends on the unknown parameter and serves as the interferometer’s output signal.
By repeating this process—preparing the system, allowing it to evolve, and measuring the population difference—a statistical dataset is collected. Techniques such as least squares method, maximum likelihood estimation or Bayesian inference can then be used to analyze the data and estimate the parameter with high precision. In the following, we will consider the estimation qualities of three differently prepared initial states: coherent spin state, squeezed spin state, and maximally entangled spin state.
@@
Figure 2: Schematic representation of the Ramsey interferometry protocol. Shown for (a) a classical (coherent) atomic ensemble and (b) a spin-squeezed atomic state. The red cloud-like shape indicates the quantum state’s uncertainty distribution, and the curved arrows denote the axes and directions of rotation. The protocol begins with an initial state polarized along the $\hat{z}-$axis. A first $\pi/2$ pulse rotates the state onto the equator of the Bloch sphere, aligning it along the $\hat{y}-$axis. During the free evolution period of duration $t$, the state precesses around the $z$-axis by an angle $\omega t$, where $\omega$ is the frequency to be estimated. A second $\pi/2$ pulse, applied around the $\hat{x}-$axis, converts the acquired phase information into a population imbalance along $\hat{z}$, which is read out via a measurement of $\langle\hat{S}_{z}\rangle$. In the spin-squeezed case (b), the reduced quantum noise along the measurement axis suppresses uncertainty in the population imbalance, enhancing the SNR and thereby improving the precision of frequency estimation. The histograms on the right side represent the probability $P(m)=|\langle m|\psi\rangle|^{2}$ of measuring population imbalance $m$. The probability distribution is narrower for a squeezed state reflecting reduced fluctuations (squeezing).
@@ IV.2 Coherent spin state
In Ramsey interferometry, states prepared in a classical superposition are called coherent spin states (CSS) [See Fig. 4(a)] and are often measured through a population difference $\sum_{i=1}^{N}\hat{\sigma}_{z}^{(i)}$. As atoms behave collectively, we define collective spin operators as
@@
which rotates around the Bloch sphere with frequency $\omega$, the expectation value of the population difference after the second $\pi/2$ pulse becomes
@@
$(\Delta\hat{S}_{z})^{2}=(\Delta\hat{S}_{y}(0))^{2}=\frac{N}{4}\;.$ (31)
@@
The maximum SNR, $S(\omega)=Nt^{2}$, defines the standard quantum limit for coherent dynamics. This is a manifestation of the central limit theorem: each of the $N$ spins behaves independently—their measurement outcomes are averaged to reduce uncertainty. In this case the intrinsic uncertainty due to quantum measurement averages out like classical noise from repeated trials. The CSS, being an unentangled product state, is relatively easy to prepare and robust to noise such as decoherence effects, making it highly attractive from a practical point of view. Atomic clocks use this principle to measure the detuning $\delta=\omega_{d}-\omega$, where $\omega_{d}$ is the laser frequency driving the system. By tuning the laser close to resonance $\cos(\delta t)\approx 1$], the SNR approaches the standard quantum limit. This enables atomic clocks to achieve extraordinary precision, detecting effects like gravitational shifts—energy level changes caused by the curvature of space-time. Over a timescale comparable to the age of the Universe, approximately 13.8 billion years, these clocks would accumulate an error of only about one second [125].
@@ IV.3 Redistribution of noise with spin-squeezing
While reaching the standard quantum limit is remarkable, further precision can be achieved by reducing quantum noise. This is done using quantum squeezing [see Fig. 4(b) and (c)], a form of entanglement where the uncertainty in one observable (for instance, $\hat{S}_{z}$) is reduced at the expense of increased uncertainty in a conjugate observable (for instance, $\hat{S}_{y}$[126]. The ultimate limit of squeezing is set by the Heisenberg uncertainty relation
@@
A natural question is how such squeezed states can be generated. The most widely studied mechanisms are one-axis twisting and two-axis counter-twisting, originally introduced by Kitagawa and Ueda [127]. In one-axis twisting, the collective spin evolves under an interaction Hamiltonian of the form $\hat{H}\propto\hat{S}_{z}^{2}$. Geometrically, this causes the initially symmetric uncertainty distribution on the Bloch sphere to twist around the $\hat{z}-$axis. As a result, the quantum noise ellipse becomes squeezed along one direction and stretched along the orthogonal direction [See Fig. 4(b)]. In contrast, two-axis counter-twisting arises from a Hamiltonian of the form $\hat{H}\propto\hat{S}_{x}^{2}-\hat{S}_{y}^{2}$. On the Bloch sphere, this corresponds to simultaneous twisting around two orthogonal axes, which redistributes quantum fluctuations more efficiently and can generate stronger squeezing than one-axis twisting [See Fig. 4(c)].
While one-axis twisting is experimentally more accessible and underlies many spin-squeezing protocols in cold atoms and trapped ions, two-axis counter-twisting represents the optimal scenario for achieving the fastest squeezing towards the Heisenberg limit [see Fig. 3]. For an optimally squeezed state generated by two-axis counter-twisting [see Fig. 4(c)], the noise reduces to
$(\Delta\hat{S}_{z}(t))^{2}\approx\frac{1}{2}\;,$ (34)
and the expense of reducing the signal by half. This overall improves the optimal SNR to
@@ IV.4 Maximally entangled states and the Heisenberg limit
Figure 3: Squeezing with two-axis counter-twisting. Improving the sensitivity with squeezing relies on reducing the noise and keeping the signal as constant as possible. Optimal state for SNR (a) is not an optimally squeezed state (b) because too much squeezing reduces the signal. There is a trade-off between these two quantitties (c). Here $N=100$.
@@
which maximizes the variance of $\hat{S}_{y}$. Depending on the field, such states are known under different names: GHZ states in the many-body and quantum information community [128], Schrödinger cat states in quantum optics [129], and NOON (N00N) states in interferometry and photonics [130, 131]. More generally, they are also described as macroscopic superposition states or maximally path-entangled states. For these states, the QFI attains its maximal value,
@@
This upper bound is four times as large as the SNR obtained from population difference measurements for optimally squeezed states, indicating that population difference measurement is not the optimal strategy. Nevertheless, population difference measurements remain near-optimal and are much simpler to implement in practice.
While maximally entangled states offer the highest QFI, their practical use is limited by significant challenges. For maximally entangled states, the population difference signal $\langle\hat{S}_{z}\rangle$ vanishes. In Ramsey interferometry with maximally entangled states, one typically relies on measuring a collective coherence $\hat{\sigma}_{x}^{\otimes N}$. In practical settings, it is typically more difficult to measure a collective coherence than a population difference (previously used for CSS and squeezed spin states). For arbitrary system size $N$ one obtains on average a measurement signal of
@@ IV.5 Limitations in open quantum systems
Thus far, our discussion has focused on closed quantum systems undergoing coherent evolution—idealized scenarios that, while insightful, are not fully representative of real-world quantum systems. In reality, no quantum system exists in complete isolation. Every atom—or any quantum system—interacts with its surrounding environment. For instance, atoms couple to the electromagnetic vacuum, which contains infinitely many modes of the electromagnetic field. This interaction leads to phenomena such as spontaneous emission, where an excited atom irreversibly emits a photon into the vacuum. Spontaneous emission fundamentally limits the lifetime of an atom’s excited state and has profound implications for Ramsey interferometry and quantum metrology whatsoever.
Figure 4: Mercator projection of the phase space picture (Husimi function) for various quantum states used in Ramsey interferometry. (a) depicts a CSS which maximally leads to standard quantum limit of precision; (b) Depicts an optimally squeezed state generated with one-axis twisting Hamiltonian which can overcome the standard quantum limit; (c) depicts an optimally squeezed state generated with two-axis counter-twisting Hamiltonian, reaching the quadratic (Heisenberg) scaling of the precision; and (d) depicts the maximally entangled state which can be used to reach the ultimate limit of precision—the Heisenberg limit. Here $N=10$.
Paradoxically, while environmental coupling is a source of decoherence and dissipation, it is also what makes quantum measurement possible. Without this interaction, quantum systems would remain entirely isolated, inaccessible to observation. The challenge in quantum metrology lies in carefully balancing the detrimental effects of decoherence with the necessity of coupling for measurement. Various models describe the interplay between quantum systems and their environments, accounting for decoherence and dissipation in different contexts [132, 133, 134, 135, 136, 137, 23, 138, 139, 133]. These effects impose constraints on the time over which a quantum system can maintain coherent evolution. For example, in the presence of dissipation, the optimal SNR no longer exhibits a simple quadratic dependence on time. Instead, it often takes the form
@@
where: $\Gamma$ quantifies the dissipation strength (for instance, the inverse lifetime of an atomic excited state); $T$ is the optimal evolution time, which often is no longer a freely tunable parameter but is determined by the interplay between quantum dynamics and noise; and $f_{N}$ is a function that depends on the number of atoms $N$, the initial state, and other parameters of the system. Importantly, $f_{N}$ is asymptotically a linear function of $N$ [137]. This relationship underscores the central role of decoherence in metrology: the maximum achievable precision is not solely dictated by quantum mechanics but also by how effectively dissipation and noise can be mitigated or leveraged. These aspects will be analyzed further in Secs. IX and X.

### v2

@@ III.7 Invertibility of the quantum Fisher information matrix
Beyond the issues discussed above, the invertibility of the Fisher information matrices plays a crucial role in multiparameter estimation. In the case of an invertible QFIM but singular (non-invertible) CFIM, the issue is associated to the incompleteness of the chosen measurement setup. Therefore, the solution is to enlarge the measurement space, for instance, by replacing projective measurements with generalized POVMs, or by employing sequential projective measurements [128, 129, 130]. On the other hand, a singular QFIM is more problematic as it directly implies that the CFIM is necessarily singular for all possible choices of the measurement setup. Since the QCRB is only well defined for an invertible QFIM, we strictly require
@@
When $\text{det}[\hat{\mathcal{I}}]=0$, the QCRB becomes undefined and nothing meaningful can be said about the achievable precision for the parameter set $\vec{\omega}$ [131, 132, 33, 133, 128, 32, 134]. A singular QFIM implies diverging variances and thus a vanishing SNR for parameter estimation. Such singularities typically arise from underlying dependencies between the original parameter set, and are such that signals corresponding to different parameters are indistinguishable in the measurement data [33]. Furthermore, recent works have demonstrated that under appropriate conditions, the CRB in the multiparameter case can be surpassed provided a probe has proximity to a singular point of the QFIM [135].
@@ III.8 From abstract to applied
To provide a practical intuition, in Section IV we revisit some of the core concepts of parameter estimation within the context of a concrete physical example: Ramsey interferometry. This paradigmatic quantum sensing protocol provides a natural setting to illustrate how quantities like the SNR, QFI, and the CRB arise in practice. We further investigate how measurement strategies and quantum resources, including entanglement and squeezing, can be exploited to enhance the sensitivity to the target parameters $\omega$. This allows us to introduce the fundamental precision limits of classical and quantum sensing: the standard quantum limit, and the Heisenberg limit. By working through this example step by step, we reinforce the abstract ideas introduced here, and lay the groundwork for more advanced schemes discussed later in this tutorial.
Building on the paradigmatic sensing strategy introduced in Section IV, we turn to the harmonic oscillator and Landau–Zener models as minimal settings to explore quantum criticality in metrology in Sections V and VI, respectively. These models allow us to highlight the features that make critical systems powerful platforms for sensing, such as their geometric structure and diverging susceptibilities. We present a detailed analysis of both single- and multiparameter estimation, focusing on the fundamental precision trade-offs, measurement incompatibility, and the emergence of singularities in the QFIM and how to deal with them. Particular focus is devoted to the role of resources, such as interrogation time, and to the impact of thermal fluctuations on critical quantum sensing.
@@ IV Parameter estimation in quantum metrology on the example of Ramsey interferometry
Importantly, the QFI can be thought of as the SNR optimized over all possible choices of measurement. However, this optimal choice may not be directly implementable. Nonetheless, it provides a crucial benchmark: by establishing the ultimate precision limit, it allows us to assess how well feasible measurement strategies perform in comparison, and to understand how close we are to the fundamental quantum bound. Consequently, a central concept in metrology and parameter estimation is the SNR—quantifying the quality of a specific measurement by comparing the strength of the signal to the noise that obscures it. With the QFI playing the role of an optimal SNR, this establishes the ultimate benchmark for precision and the link between abstract theoretical limits to real-world measurement challenges.
@@ IV.1 Ramsey interferometry
To illustrate the principles of quantum metrology, we consider the Ramsey interferometer—a cornerstone of precision measurement and the operating principle behind atomic clocks, which rank among the most accurate instruments ever built [136]. Their extraordinary precision stems from the inherent stability of atomic transition frequencies: each atom functions as an almost perfect clock, and its ticks are dictated by fundamental physical laws, much like those of an ideal pendulum. By counting the oscillations of an atom’s quantum state, time can be tracked with remarkable accuracy. In practice, however, imperfections such as fluctuations in the probing laser’s intensity and frequency introduce unavoidable sources of imprecision.
@@
Figure 2: Schematic representation of the Ramsey interferometry protocol. The protocol is illustrated for (a) a classical (coherent) atomic ensemble and (b) a spin-squeezed atomic state. The red cloud-like shape indicates the quantum state’s uncertainty distribution, and the curved arrows denote the axes and directions of rotation. The protocol begins with an initial state polarized along the $z$-axis. A first $\pi/2$ pulse rotates the state onto the equator of the Bloch sphere, aligning it along the $y$-axis. During the free evolution period of duration $t$, the state precesses around the $z$-axis by an angle $\omega t$, where $\omega$ is the frequency to be estimated. A second $\pi/2$ pulse, applied around the $x$-axis, converts the acquired phase information into a population imbalance along $z$, which is read out via a measurement of $\langle\hat{S}_{z}\rangle$. In the spin-squeezed case (b), the reduced quantum noise along the measurement axis suppresses uncertainty in the population imbalance, enhancing the SNR and thereby improving the precision of frequency estimation. The histograms on the right side represent the probability $P(m)=|\langle m|\psi\rangle|^{2}$ of measuring population imbalance $m$. The probability distribution is narrower for a squeezed state reflecting reduced fluctuations (squeezing).
b
In a Ramsey interferometer [see Fig. 2 for a schematic Bloch sphere representation], parameter estimation begins by preparing a quantum system in a well-defined initial state, typically the collective ground state. The first step is to apply a pulse that creates a coherent superposition of the two energy levels, represented as a quantum state vector on the Bloch sphere, halfway between the two energy eigenstates.
Next, the system evolves freely under the influence of the unknown parameter, such as a frequency, magnetic or gravitational field. During this evolution, a relative phase accumulates between the components of the superposition state, encoding information about the parameter. The evolution time is carefully chosen to maximize sensitivity to small variations in the parameter. Because directly measuring the phase is challenging, a second pulse is applied after the evolution period. This pulse converts the accumulated phase information into a measurable quantity, typically the population difference between the two energy levels. The population difference depends on the unknown parameter, and serves as the interferometer’s output signal.
By repeating this process—preparing the system, allowing it to evolve, and measuring the population difference—a statistical dataset is collected. Techniques such as least squares method, maximum likelihood estimation, or Bayesian inference can then be used to analyze the data and estimate the parameter with high precision. In the following, we will consider the estimation qualities of three differently prepared initial states: coherent spin state, squeezed spin state, and maximally entangled spin state.
@@ IV.2 Coherent spin state
In Ramsey interferometry, states prepared in a classical superposition are called coherent spin states (CSS) [137] [see Fig. 3(a)], and are often measured through a population difference $\sum_{i=1}^{N}\hat{\sigma}_{z}^{(i)}$. As atoms behave collectively, we define collective spin operators as
@@
which rotates around the Bloch sphere with frequency $\omega$, the expectation value of the population difference after the second $\pi/2$ pulse is
@@
$(\Delta\hat{S}_{z})^{2}=[\Delta\hat{S}_{y}(t=0)]^{2}=\frac{N}{4}\;.$ (31)
@@
The maximum SNR, $S_{\omega}=Nt^{2}$}, defines the standard quantum limit for coherent dynamics. This is a manifestation of the central limit theorem: each of the $N$ spins behaves independently—their measurement outcomes are averaged to reduce uncertainty. In this case the intrinsic uncertainty due to quantum measurement averages out like classical noise from repeated trials. The CSS, being an unentangled product state, is relatively easy to prepare and robust to noise such as decoherence effects, making it highly attractive from a practical point of view. Atomic clocks use this principle to measure the detuning $\delta=\omega_{d}-\omega$}, where $\omega_{d}$ is the laser frequency driving the system. By tuning the laser close to resonance $\cos(\delta t)\approx 1$}], the SNR approaches the standard quantum limit. This enables atomic clocks to achieve extraordinary precision, detecting effects like gravitational shifts—energy level changes caused by the curvature of space-time. Over a timescale comparable to the age of the Universe, approximately 13.8 billion years, these clocks would accumulate an error of only about one second [138].
@@ IV.3 Redistribution of noise with spin-squeezing
While reaching the standard quantum limit is remarkable, further precision can be achieved by reducing quantum noise. This is done using quantum squeezing [see Fig. 3(b) and (c)], a form of entanglement where the uncertainty in one observable (for instance $\hat{S}_{z}$) is reduced at the expense of increased uncertainty in a conjugate observable (for instance $\hat{S}_{y}$[139]. The ultimate limit of squeezing is set by the Heisenberg uncertainty relation
@@
A natural question is how such squeezed states can be generated. The most widely studied mechanisms are one-axis twisting and two-axis counter-twisting, originally introduced by Kitagawa and Ueda [140]. In one-axis twisting, the collective spin evolves under an interaction Hamiltonian of the form $\hat{H}\propto\hat{S}_{z}^{2}$. Geometrically, this causes the initially symmetric uncertainty distribution on the Bloch sphere to twist around the $z$-axis. As a result, the quantum noise ellipse becomes squeezed along one direction and stretched along the orthogonal direction [see Fig. 3(b)]. In contrast, two-axis counter-twisting arises from a Hamiltonian of the form $\hat{H}\propto\hat{S}_{x}^{2}-\hat{S}_{y}^{2}$. On the Bloch sphere, this corresponds to simultaneous twisting around two orthogonal axes, which redistributes quantum fluctuations more efficiently and can generate stronger squeezing than one-axis twisting [see Fig. 3(c)].
While one-axis twisting is experimentally more accessible, and underlies many spin-squeezing protocols in cold atoms and trapped ions, two-axis counter-twisting represents the optimal scenario for achieving the fastest squeezing towards the Heisenberg limit [see Fig. 4]. For an optimally squeezed state generated by two-axis counter-twisting [see Fig. 3(c)], the noise reduces to
$[\Delta\hat{S}_{z}(t)]^{2}\approx\frac{1}{2}\;,$ (34)
where this noise reduction comes at the expense of reducing the signal by half. This overall improves the optimal SNR to
@@
Figure 3: Mercator-projected phase-space (Husimi Q) distributions of quantum states used in Ramsey interferometry. Shown are: (a) CSS achieving the standard quantum limit; (b) optimally squeezed state generated via a one-axis twisting Hamiltonian, capable of surpassing the standard quantum limit; (c) optimally squeezed state generated via a two-axis counter-twisting Hamiltonian, reaching the quadratic (Heisenberg) scaling; (d) maximally entangled state achieving the ultimate Heisenberg limit. To obtain the plots, we set $N=10$.
@@ IV.4 Maximally entangled states and the Heisenberg limit
which maximizes the variance of $\hat{S}_{y}$. Depending on the field, such states are known under different names: GHZ states in the many-body and quantum information community [141], Schrödinger cat states in quantum optics [142], and NOON (N00N) states in interferometry and photonics [143, 144]. More generally, they are also described as macroscopic superposition states or maximally path-entangled states. For these states, the QFI attains its maximal value
@@
This upper bound is four times as large as the SNR obtained from population difference measurements for optimally squeezed states, indicating that population difference measurement is not the optimal strategy. Nevertheless, population difference measurements remain near-optimal, and are much simpler to implement in practice.
While maximally entangled states offer the highest QFI, their practical use is limited by significant challenges. For these states, the population difference signal $\langle\hat{S}_{z}\rangle$ vanishes. In Ramsey interferometry with maximally entangled states, one typically relies on measuring a collective coherence $\hat{\sigma}_{x}^{\otimes N}$. In practical settings, it is typically more difficult to measure a collective coherence than a population difference (previously used for CSS and squeezed spin states). For arbitrary system size $N$ one obtains on average a measurement signal of
@@
Figure 4: Squeezing with two-axis counter-twisting. Mercator-projected phase-space (Husimi Q) distributions of: (a) the state that maximizes SNR, and (b) the state with minimal quantum fluctuations (maximally squeezed). We show that the state maximizing the SNR is not maximally squeezed, because excessive squeezing reduces the signal. Panel (c) illustrates the resulting trade-off between signal strength and noise reduction. To obtain the plots, we set $N=100$ and evolved an initial coherent state polarized along $x$ axis with a two-axis counter-twisting unitary $\hat{U}(t)=\exp[-i\chi t(\hat{S}_{z}^{2}-\hat{S}_{y}^{2})]$.
@@ IV.5 Limitations in open quantum systems
Thus far, our discussion has focused on closed quantum systems undergoing coherent evolution—idealized scenarios that, while insightful, are not fully representative of real-world quantum systems. In reality, no quantum system exists in complete isolation. Every atom—or any quantum system—interacts with its surrounding environment. For instance, atoms couple to the electromagnetic vacuum, which contains infinitely many modes of the electromagnetic field. This interaction leads to phenomena such as spontaneous emission, where an excited atom irreversibly emits a photon into the vacuum. Spontaneous emission fundamentally limits the lifetime of an atom’s excited state, and has profound implications for Ramsey interferometry and quantum metrology whatsoever.
Paradoxically, while environmental coupling is a source of decoherence and dissipation, it is also what makes quantum measurement possible. Without this interaction, quantum systems would remain entirely isolated, inaccessible to observation. The challenge in quantum metrology lies in carefully balancing the detrimental effects of decoherence with the necessity of coupling for measurement. Various models describe the interplay between quantum systems and their environments, accounting for decoherence and dissipation in different contexts [145, 146, 147, 148, 149, 150, 25, 151, 152, 146]. These effects impose constraints on the time over which a quantum system can maintain coherent evolution. For example, in the presence of dissipation the optimal SNR no longer exhibits a simple quadratic dependence on time. Instead, it often takes the form
@@
where: $\Gamma$ quantifies the dissipation strength (for instance, the inverse lifetime of an atomic excited state); $T$ is the optimal evolution time, which often is no longer a freely tunable parameter but is determined by the interplay between quantum dynamics and noise; and $f_{N}$ is a function that depends on the number of atoms $N$, the initial state, and other parameters of the system. Importantly, $f_{N}$ is asymptotically a linear function of $N$ [150]. This relationship underscores the central role of decoherence in metrology: the maximum achievable precision is not solely dictated by quantum mechanics but also by how effectively dissipation and noise can be mitigated or leveraged. These aspects will be analyzed further in Sections IX and X.